# Global self--similarity of dense granular flow in hopper: the role of hopper width

### v2

@@ Abstract
keywords:
Granular materials , Dense granular flow , Flow rate , Hopper width , Discrete element method
journal: XXX
1 Introduction
@@
Beverloo law is a classical empirical formula describing the relationship between the dense flow rate $Q$ and the outlet size $D$ by using a simple physical analysis.[17]
@@
In this study, both experiments and numerical simulations are conducted to investigate the impact of hopper width on dense granular flow in a two--dimensional hopper. The experimental setup and corresponding simulation model are detailed in Section 2. In Section 3, we first explore the influence of hopper width on the dense granular flow rate, followed by an examination of the relationship between the dense flow rate and outlet size using the Beverloo and Janda laws. Subsequently, we analyze the flow properties at both local and global scales. We revisit the self--similarities concerning packing density and grain velocity around the outlet and introduce the global similarity regarding the transition between the mass flow pattern and the funnel flow pattern. Finally, our main results are summarized in Section 4.
2 Experiment setup and simulation model
@@
Fig. 1: Snapshots of (a) experiment setup and (b) simulation system.
@@
3 Results and discussion
@@
Fig. 2: (a) Outflow rate as a function of hopper width for experiment and simulation results. The outlet size is fixed at $D=18~{}{\rm mm}$. The dash and the dotted lines donate the mean flow rate of large hopper widths. (b) Outflow rate normalized by the mean flow rate of large hopper widths as a function of scaled baffle length. The dash and the dotted lines donate the fitted results for $D/d=9$ by using Eq.(5). The other outlet sizes, $D/d=7,10,11$ and 12, are used in the simulations and donated by different types of symbols.
@@
Fig. 3: Simulated results of flow rate as a function of outlet size for different baffle lengths $W_{\rm B}=1.5$ and $15d$. The red dash curves and blue dotted curves are obtained by using the Beverloo and the Janda equations, Eqs.(1) and (2), respectively.
@@
Fig. 4: Spatial distribution of (a)-(d) contact force between grains and (a)-(d) the kinetic force. Four baffle lengths are used, i.e., (a)(a)$W_{\rm B}/d=22.5$, (b)(b)$W_{\rm B}/d=15$, (c)(c)$W_{\rm B}/d=5$, (d)(d)$W_{\rm B}/d=1.5$.
@@
Fig. 5: Spatial distribution of (a)-(d) packing density and (a)-(d) grain velocity for different baffle lengths. The outlet size is $D=18~{}{\rm mm}$. The solid lines of different colors indicate the contours of different values.
@@
Fig. 6: Self--similar horizontal profiles of (a) vertical velocity, (b) packing density, and (c) flow rate, and corresponding normalized profiles of (a) vertical velocity, (b) packing density, and (c) flow rate for different baffle widths. $x_{\rm c}$ is the location of outlet center. The outlet size is $D=18~{}{\rm mm}$.
@@
Fig. 7: Packing density (squares) and mean vertical velocity (circles) at the center of the exit as a function of baffle length. The width of the outlet is $D=18~{}{\rm mm}$. The dashed curve is obtained by using Eq.(6).
@@
Fig. 8: (a) Snapshot of extraction area for granular vertical velocity. Vertical grain velocity of selected areas as a function of the packing height for baffle length (b) $W_{\rm B}/d=22.5$, (c) $W_{\rm B}/d=15$, (d) $W_{\rm B}/d=5$ and (e) $W_{\rm B}/d=1.5$. The width of the outlet is $D=18~{}{\rm mm}$.
Fig. 9: Relationship between the transition height and the baffle length. Circles are the simulation results. Dashed line is the fitting result by using the equation $H_{\rm Tr}=H_{\rm Tr}^{0}[1+{\lambda_{1}}e^{-\frac{W_{\rm B}}{{\lambda_{2}}d}}]$. Dotted line indicates the critical transition height ${H_{\rm Tr}^{0}}=28$.
@@ 4 Conclusions
The variance in baffle length has little influence on grain velocity in the local regions around the outlet; however, a similar exponential change, like the relationship between dense flow rate and baffle length, is observed for the first time. Furthermore, the dependence of the transition height from the mass flow pattern to the funnel flow pattern on the baffle length is also identified for the first time. These results demonstrate a global similarity of the flow behaviors throughout the entire hopper in terms of exponents. The earlier occurrence of the transition from the mass flow pattern to the funnel flow pattern leads to higher grain velocity, resulting in a larger flow rate. These findings indicate that a similar transition between the interior flow patterns near and far from the outlet affects the dense flow rate in the hopper. This exponential similarity also provides insights for developing a general constitutive framework for dense granular flow.

### v3

@@ Abstract
pacs:
45.70.Mg, 64.60.-i, 75.40.Gb
I Introduction
@@
In classical research, differential relationship between the physical quantities is the key point to define the dynamical characteristics. Given this consideration, Beverloo law is a successful empirical formula describing the relationship between the dense flow rate $Q$ and the outlet size $D$ by using a simple physical analysis.[17]
@@
In this study, both experiments and numerical simulations are conducted to investigate the impact of hopper width on dense granular flow in a two--dimensional hopper. The experimental setup and corresponding simulation model are detailed in Section II. In Section III, we first explore the influence of hopper width on the dense granular flow rate, followed by an examination of the relationship between the dense flow rate and outlet size using the Beverloo and Janda laws. Subsequently, we analyze the flow properties at both local and global scales. We revisit the self--similarities concerning packing density and grain velocity around the outlet and introduce the global similarity regarding the transition between the mass flow pattern and the funnel flow pattern. Finally, our main results are summarized in Section IV.
II Experiment setup and simulation model
@@
Figure 1: (Color online). Snapshots of (a) experiment setup and (b) simulation system.
@@
Figure 2: (Color online). (a) Outflow rate as a function of hopper width for experiment and simulation results. The outlet size is fixed at $D=18~{}{\rm mm}$. The dash and the dotted lines donate the mean flow rate of large hopper widths. (b) Outflow rate normalized by the mean flow rate of large hopper widths as a function of scaled baffle length. The dash and the dotted lines donate the fitted results for $D/d=9$ by using Eq.(5). The other outlet sizes, $D/d=7,10,11$ and 12, are used in the simulations and donated by different types of symbols.
III Results and discussion
@@
Figure 3: (Color online). Simulated results of flow rate as a function of outlet size for different baffle lengths $W_{\rm B}=1.5$ and $15d$. The red dash curves and blue dotted curves are obtained by using the Beverloo and the Janda equations, Eqs.(1) and (2), respectively.
@@
Figure 4: (Color online). Spatial distribution of (a)-(d) contact force between grains and (a)-(d) the kinetic force. Four baffle lengths are used, i.e., (a)(a)$W_{\rm B}/d=22.5$, (b)(b)$W_{\rm B}/d=15$, (c)(c)$W_{\rm B}/d=5$, (d)(d)$W_{\rm B}/d=1.5$.
@@
Figure 5: (Color online). Spatial distribution of (a)-(d) packing density and (a)-(d) grain velocity for different baffle lengths. The outlet size is $D=18~{}{\rm mm}$. The solid lines of different colors indicate the contours of different values.
@@
Figure 6: (Color online). Self--similar horizontal profiles of (a) vertical velocity, (b) packing density, and (c) flow rate, and corresponding normalized profiles of (a) vertical velocity, (b) packing density, and (c) flow rate for different baffle widths. $x_{\rm c}$ is the location of outlet center. The outlet size is $D=18~{}{\rm mm}$.
@@
Figure 7: (Color online). Packing density (squares) and mean vertical velocity (circles) at the center of the exit as a function of baffle length. The width of the outlet is $D=18~{}{\rm mm}$. The dashed curve is obtained by using Eq.(6).
@@
Figure 8: (Color online). (a) Snapshot of extraction area for granular vertical velocity. Vertical grain velocity of selected areas as a function of the packing height for baffle length (b) $W_{\rm B}/d=22.5$, (c) $W_{\rm B}/d=15$, (d) $W_{\rm B}/d=5$ and (e) $W_{\rm B}/d=1.5$. The width of the outlet is $D=18~{}{\rm mm}$.
Figure 9: Relationship between the transition height and the baffle length. Circles are the simulation results. Dashed line is the fitting result by using the equation $H_{\rm Tr}=H_{\rm Tr}^{0}[1+{\lambda_{1}}e^{-\frac{W_{\rm B}}{{\lambda_{2}}d}}]$. Dotted line indicates the critical transition height ${H_{\rm Tr}^{0}}=28$.
@@ IV Conclusions
The variance in baffle length has little influence on grain velocity in the local regions around the outlet; however, a similar exponential change, like the relationship between dense flow rate and baffle length, is observed for the first time. Furthermore, the dependence of the transition height from the mass flow pattern to the funnel flow pattern on the baffle length is also identified for the first time. These results demonstrate a global similarity of the flow behaviors throughout the entire hopper in terms of exponents. The earlier occurrence of the transition from the mass flow pattern to the funnel flow pattern leads to higher grain velocity, resulting in a larger flow rate. These findings indicate that a similar transition between the interior flow patterns near and far from the outlet affects the dense flow rate in the hopper. This exponential similarity also provides further insights for developing a general constitutive framework for dense granular flow.